how many cubes, with side measures of 2 cm, will fit inside a right rectangular prism with dimensions of 6 cm by 8 cm by 4 1/2 cm?
Answer by stanbon(75887)

how many cubes, with side measures of 2 cm, will fit inside a right rectangular prism with dimensions of 6 cm by 8 cm by 4 1/2 cm?
--------
Volume of each cube = 2^3 = 8cm^3
--------
Volume of the prism = 6*8*4.5 = 216 cm^3
-----
# of cubes = 216/8 = 27
===========================
Cheers,
Stan H.
---------------------------